40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Thu Dec 18 02:23:34 2014, ericq, Update, LSC, Locking update 15lockloss.png15spec.pdf
    Reply  Fri Dec 19 15:59:49 2014, Jenne, Update, LSC, Lockloss from Wed 1102929772_PCdriveRailed.png
Message ID: 10814     Entry time: Thu Dec 18 02:23:34 2014     Reply to this: 10818
Author: ericq 
Type: Update 
Category: LSC 
Subject: Locking update 

 [ericq, Diego]

Some locking efforts tonight; many locklosses due to PRC angular motion. Furthest progress was arm powers of 15, and I've stared at the corresponding lockloss plot, with little insight into what went wrong. (BTW, lastlock.sh seems to catch the lock loss reliably in the window)

15lockloss.png

CARM and DARM loops were measured not long before this lock loss, and had nominal UGFs (~120Hz, ~20deg PM). However, there was a reasonably clear 01 mode shape at the AS camera, which I did nothing to correct. Here's a spectrum from *just* before the lockloss, recovered via nds. Nothing stands out to me, other than a possible loss of DARM optical gain. (I believe the references are the error signal spectra taken in ALS arms held away + PRMI on 3F configuration)

15spec.pdf

The shape in the DARM OLTF that we had previously observed and hypothesized as possible DARM optical spring was not ever observed tonight. I didn't induce a DARM offset to try and look for it either, though.

Looking into some of the times when I was measuring OLTFs, the AS55 signals do show coherence with the live DARM error signal at the excitation frequencies, but little to no coherence under 30Hz, which probably means we weren't close enough to swap DARM error signals yet. This arm power regime is where the AS55 sign flip has been modeled to be...


A fair amount of time was spent in pre-locking prep, including:

  • The usual X green beat alignment tweak, to make the X ALS control not terrible
  • Both ITM oplevs were centered
  • For some reason, I had to flip the signs of the REFL165 matrix elements for the PRMI...
  • "Restore PRMI SB" has ASC automatically enabled, which results in unexpected kicks even with LSC off, which caused a few minutes head-scratching
ELOG V3.1.3-